if x³+3x²+3x+1 is divided by x-½ then the remainder
Answer by stanbon(75887)

if x³+3x²+3x+1 is divided by x-½ then the remainder
is f(1/2) = (1/2)^3 + 3(1/2)^2 + 2(1/2) + 1
---
= (1/8) + 3/4 + 1 + 1
---
= 2 7/8
==============
Cheers,
Stan H.
----------------